# Rendering-Oriented 3D Point Cloud Attribute Compression using Sparse Tensor-based Transformer

### v2

@@ Abstract
The evolution of 3D visualization techniques has fundamentally transformed how we interact with digital content. At the forefront of this change is point cloud technology, offering an immersive experience that surpasses traditional 2D representations. However, the massive data size of point clouds presents significant challenges in data compression. Current methods for lossy point cloud attribute compression (PCAC) generally focus on reconstructing the original point clouds with minimal error. However, for point cloud visualization scenarios, the reconstructed point clouds with distortion still need to undergo a complex rendering process, which affects the final user-perceived quality. In this paper, we propose an end-to-end deep learning framework that seamlessly integrates PCAC with differentiable rendering, denoted as rendering-oriented PCAC (RO-PCAC), directly targeting the quality of rendered multiview images for viewing. In a differentiable manner, the impact of the rendering process on the reconstructed point clouds is taken into account. Moreover, we characterize point clouds as sparse tensors and propose a sparse tensor-based transformer, called SP-Trans. By aligning with the local density of the point cloud and utilizing an enhanced local attention mechanism, SP-Trans captures the intricate relationships within the point cloud, further improving feature analysis and synthesis within the framework. Extensive experiments demonstrate that the proposed RO-PCAC achieves state-of-the-art compression performance, compared to existing reconstruction-oriented methods, including traditional, learning-based, and hybrid methods.
@@ I Introduction
    we make the first attempt to construct a general rendering-oriented compression framework for the color attribute of 3D point cloud data, termed RO-PCAC. RO-PCAC consists of a learning-based point cloud attribute compression module and a differentiable rendering module, enabling end-to-end training to jointly minimize the bitrate and optimize the quality of rendered multiview images. Furthermore, RO-PCAC supports the modular design for both the compression and rendering components.
@@ II-A Point Cloud Geometry Compression
Traditional point cloud geometry coding has mature frameworks, mainly dominated by MPEG G-PCC and V-PCC standards [10, 11, 12]. These frameworks have set the benchmark for efficient compression techniques. The MPEG G-PCC codec, for instance, employs an octree-based partitioning strategy to hierarchically structure point clouds [34], which allows for efficient compression by encoding the octree’s occupancy. G-PCC also incorporates a surface modeling technique [10] for point clouds with dense sampling. Surface approximation is achieved through the use of triangular meshes, or ’trisoup’, which is handled by a dedicated geometry encoding scheme. In a divergent methodology, V-PCC begins the compression process by converting the data from three-dimensional to two-dimensional space, enabling the use of established video encoding standards such as High-Efficiency Video Coding (HEVC) [35] for encoding planes and depth information across a sequence. Currently, G-PCC currently stands as a leading protocol in terms of compression efficiency for static point cloud data.
@@ II-C Learning-based Point Cloud Attribute Compression
The success of sparse convolution in compressing point cloud geometry has led to its application for attribute compression as well. Wang et al. modeled the point clouds as sparse tensors and processed them by a series of stacked sparse convolution layers, termed SparsePCAC, which significantly narrowed the performance gap with the traditional G-PCC [21]. Zhang et al. presented ScalablePCAC, a scalable compression technique for point cloud attributes [22]. ScalablePCAC integrates a standard G-PCC codec as the base layer and a sparse convolution-based model as the enhancement layer to handle full-resolution point cloud compression. An investigation of different quantization parameter combinations of the two layers is implemented to improve the rate allocation. Further, Figueiredo et al. proposed an inherited embedded attribute encoding method by set partitioning in hierarchical trees and an MLP for context modeling [23]. Full scalability is achieved with a single code stream. In this paper, we focus on exploring the potential of end-to-end deep learning-based point cloud compression methods. Consequently, we compare the performance of our approach with the latest traditional point cloud compression method (i.e., G-PCC), the state-of-the-art deep learning-based methods (i.e., SparsePCAC) and a hybrid approach of traditional method and deep learning-based method (i.e., ScalablePCAC).
@@ III-A3 Overview
Fig. 2 illustrates the flowchart of the proposed rendering-oriented point cloud compression framework. The framework is divided into two primary modules: the compression module and the rendering module. The compression module efficiently compresses the input point clouds to compact code streams and decompresses them to reconstructed point clouds. The rendering module renders the reconstructed point clouds to images from preset multiple perspectives through differentiable rendering techniques. Note that the rendering component of our framework is modular, allowing for the potential integration of future advancements in differentiable rendering techniques to replace the current rendering module used in this study.
@@ III-B Compression Module
where $\vec{f}_{i}^{\prime}$ is the output feature and $\delta_{\mathrm{rel}}\left(\vec{c}_{i}-\vec{c}_{j}\right)$ is the relative position encoding. Functions $\theta$, $\alpha$ and $\lambda$ are the multi-layer perceptron-based query, key and value projection layers, respectively. The local self-attention layer utilizes the cosine similarity to measure the relationship between the $\theta\left(\vec{f}_{i}\right)$ and $\alpha\left(\vec{f}_{j}\right)+\delta_{\mathrm{rel}}\left(\vec{c}_{i}-\vec{c}_{j}\right)$ vectors, rather than the commonly used $\operatorname{softmax}(\cdot)$ function. This choice stems from the adaptive nature of the fixed-size window. Firstly, by focusing on the direction, it inherently handles similarity and makes a fair comparison when the number of neighbors varies. In contrast, the similarity calculated from the $\operatorname{softmax}(\cdot)$ function highly depends on the number of neighbours. Secondly, it prevents the potential issue of the $\operatorname{softmax}(\cdot)$ function. A single neighbor (i.e., $|\mathcal{N}(i)|=1$) would make the attention mechanism degenerate to a linear layer by normalizing the weight to unity. Thus, using cosine similarity in the local self-attention layer of SP-Trans is a proper choice that aligns with the adaptive, fixed-size window approach to neighbor searching. It ensures equitable weighting of neighbors that reflects the local point distribution. Combined with a skip connection, the final SP-Trans is illustrated in Fig. 5.
@@ III-C Rendering Module
Once the points are in the camera view coordinate system, they need to be transformed into the normalized device coordinate (NDC) $X_{\text{ndc}}$. The specific type of camera (perspective or orthographic) will determine the projection matrix used for this transformation. The perspective camera, which more closely conforms to real scene shooting is chosen in our framework. The coordinates are first placed into a symmetric viewing frustum as shown in Fig. 7(a) and the corresponding projective transformation transform the viewing frustum into a cubic space as shown in Fig. 7(b). The corresponding projective transformation, associated with a homogeneous division can be represented as
@@
$\tilde{X}_{\text{ndc}}$ $=\left[\begin{matrix}\frac{1}{\tan\left(\frac{fov}{2}\right)}&0&0&0\\ 0&\frac{1}{\tan\left(\frac{fov}{2}\right)\times ap}&0&0\\ 0&0&\frac{far}{far-near}&-\frac{far\times near}{far-near}\\ 0&0&1&0\end{matrix}\right]\tilde{X}_{\text{view}}$ (8)
@@
$\simeq\left[\begin{matrix}\frac{x}{z\times\tan\left(\frac{fov}{2}\right)}\\ \frac{y}{z\times\tan\left(\frac{fov}{2}\right)\times ap}\\ \frac{far\left(z-near\right)}{z\left(far-near\right)}\\ 1\end{matrix}\right],$
@@
where $fov$ is the field of view angle, $near$ is the distance of the camera to the near clipping plane, $far$ is the distance to the far clipping plane and $ap$ is the aspect ratio of the image pixels. $\sim$ means the homogeneous coordinates of corresponding terms, which are used in this calculation in the projective space. $\left(x,y,z,1\right)^{\top}$ is the homogeneous coordinates of $\tilde{X}_{\text{view}}$. $\simeq$ denotes the equality under the homogeneous coordinates. From the coordinates of $\tilde{X}_{\text{ndc}}$, we can see that the horizontal and vertical coordinates are scaled by the depth value. As a result, objects that are closer to the viewer appear larger than those that are farther away. This is how the projective transformation altered the geometry of the point cloud. As shown in Fig. 8, square grids of the chessboard will appear to converge and become trapezoidal in the distance. Without loss of generality, the farther an object is from the viewer, the more distorted its shape becomes. Finally, the viewport transformation, determined by the parameters of the rendering window such as its resolution, maps the NDC coordinates, which range from $-1$ to $1$, to screen coordinates. To provide a sense of depth and realism to the rendered 2D image, these sequential transformations present these points in 3D world with a new geometric relationship in 2D image.
@@
Subsequently, based on the altered geometry, the rasterization algorithm sorts the points in descending order of their depth values $z$ to ensure that the closest points are processed first as shown in Fig. 9. This is achieved by calculating the depth value for each point and storing it in a z-buffer. Each point is splatted to a circular region in screen space, with opacity decreases away from the region’s center. The value of each pixel is computed by blending information from the K-nearest points in the z-buffer whose splatted regions overlap the pixel.
@@ IV-A1 Datasets
  • Training datasets. The model is trained on 548 point clouds sampled from 3D models in the real-world textured things (RWTT) dataset [63]. For each point cloud, $1\times 10^{6}$ points are randomly sampled from the surface and then voxelized to a $1024^{3}$ space to maintain consistency with the test datasets. We show four training samples in Fig 10(a).
  • Testing datasets. Eight point clouds from two datasets used in MPEG are included in our testing samples. As shown in Fig. 10(b) and Fig. 10(c), four point clouds are defined in 8iVFB [32] and the other four are defined in Owlii [33]. The training and testing datasets are noticeably different, which demonstrates the generalization of our framework.
@@ IV-A2 Training and testing details
  • Training phase. The ground truth multiviews include 8 views with elevations of $0^{\circ}$ and azimuths from $0^{\circ}$ to $300^{\circ}$ with an interval of $60^{\circ}$ and 2 views with elevations of $90^{\circ}$ and $270^{\circ}$ and azimuth of $0^{\circ}$, as shown in Fig. 11. The resolution of each view’s image is 384$\times$384. For the loss function, the value of $\lambda$ is set to 25,000, 4,000, 800, 250, and 85 for the training of the respective models. Each model undergoes training consisting of 140 epochs, facilitated by the PyTorch platform and the Adam optimizer. The parameters $\beta_{1}$ and $\beta_{2}$ of the Adam optimizer are configured at 0.9 and 0.999, respectively. The learning rate starts at 1e-4 and is progressively reduced to 1e-6, with a reduction to three-fourths of its previous value occurring every 15 epochs. The batch size is set to 8. The experimental setup includes a computer equipped with an Intel i9-14900k CPU operating at 3.60 GHz, 32 GB of main memory, and an NVIDIA GeForce RTX 4090 GPU.
  • Testing phase. To demonstrate the generalization of RO-PCAC to different views, the ground truth multiviews include 6 views with the elevation of $0^{\circ}$ and azimuths from $30^{\circ}$ to $330^{\circ}$ with an interval of $60^{\circ}$. The resolution of each view’s image is 1024$\times$1024. The R-D performance includes bpp for rate, PSNR and MS-SSIM for distortion. Furthermore, the BD-BR (Bjøntegaard Delta Rate) [64] is used to evaluate the overall compression efficiency. The R-D performance in both Y and YUV (6:1:1) spaces [65] is reported.

### v3

@@ Abstract
The evolution of 3D visualization techniques has fundamentally transformed how we interact with digital content. At the forefront of this change is point cloud technology, offering an immersive experience that surpasses traditional 2D representations. However, the massive data size of point clouds presents significant challenges in data compression. Current methods for lossy point cloud attribute compression (PCAC) generally focus on reconstructing the original point clouds with minimal error. However, for point cloud visualization scenarios, the reconstructed point clouds with distortion still need to undergo a complex rendering process, which affects the final user-perceived quality. In this paper, we propose an end-to-end deep learning framework that seamlessly integrates PCAC with differentiable rendering, denoted as rendering-oriented PCAC (RO-PCAC), directly targeting the quality of rendered multiview images for viewing. In a differentiable manner, the impact of the rendering process on the reconstructed point clouds is taken into account. Moreover, we characterize point clouds as sparse tensors and propose a sparse tensor-based transformer, called SP-Trans. By aligning with the local density of the point cloud and utilizing an enhanced local attention mechanism, SP-Trans captures the intricate relationships within the point cloud, further improving feature analysis and synthesis within the framework. Extensive experiments demonstrate that the proposed RO-PCAC achieves state-of-the-art compression performance, compared to existing reconstruction-oriented methods, including traditional, learning-based, and hybrid methods. The code will be released at https://github.com/net-F/RO-PCAC.git.
@@ I Introduction
    we make the first attempt to construct a generic rendering-oriented compression framework for the color attribute of 3D point cloud data, termed RO-PCAC. RO-PCAC consists of a learning-based point cloud attribute compression module and a differentiable rendering module, enabling end-to-end training to jointly minimize the bitrate and optimize the quality of rendered multiview images. Furthermore, RO-PCAC supports the modular design for both the compression and rendering components.
@@ II-A Point Cloud Geometry Compression
Traditional point cloud geometry coding has mature frameworks [4, 5, 7, 8, 9, 10, 11, 12], mainly dominated by MPEG G-PCC and V-PCC standards [10, 11, 12]. These frameworks have set the benchmark for efficient compression techniques. The MPEG G-PCC codec, for instance, employs an octree-based partitioning strategy to hierarchically structure point clouds [34], which allows for efficient compression by encoding the octree’s occupancy. G-PCC also incorporates a surface modeling technique [10] for point clouds with dense sampling. Surface approximation is achieved through the use of triangular meshes, or ’trisoup’, which is handled by a dedicated geometry encoding scheme. In a divergent methodology, V-PCC begins the compression process by converting the data from three-dimensional to two-dimensional space, enabling the use of established video encoding standards such as High-Efficiency Video Coding (HEVC) [35] for encoding planes and depth information across a sequence. Currently, G-PCC currently stands as a leading protocol in terms of compression efficiency for static point cloud data.
@@ II-C Learning-based Point Cloud Attribute Compression
The success of sparse convolution in compressing point cloud geometry has led to its application for attribute compression as well. Wang et al. modeled the point clouds as sparse tensors and processed them by a series of stacked sparse convolution layers, termed SparsePCAC, which significantly narrowed the performance gap with the traditional G-PCC [21]. Zhang et al. presented ScalablePCAC, a scalable compression technique for point cloud attributes [22]. ScalablePCAC integrates a standard G-PCC codec as the base layer and a sparse convolution-based model as the enhancement layer to handle full-resolution point cloud compression. An investigation of different quantization parameter combinations of the two layers is implemented to improve the rate allocation. Further, Figueiredo et al. proposed an inherited embedded attribute encoding method by set partitioning in hierarchical trees and an MLP for context modeling [23]. Full scalability is achieved with a single code stream. Guo et al. proposed TSC-PCAC, an end-to-end voxel transformer and sparse convolution module (TSCM) based point cloud attribute compression [45]. TSCM extracts global and local inter-point relevance to enhance feature representation. Recently, Joint Photographic Experts Group (JPEG) published the deep-learning based coding architecture, modules, and the deep learning models involved in the encoding and decoding processes, which covers various coding configurations to achieve target compression rates.[46] In this paper, we focus on exploring the potential of end-to-end deep learning-based point cloud compression methods. Consequently, we compare the performance of our approach with the latest traditional point cloud compression method (i.e., G-PCC), the state-of-the-art deep learning-based methods (i.e., SparsePCAC) and a hybrid approach of traditional method and deep learning-based method (i.e., ScalablePCAC).
@@ III-A3 Overview
Fig. 2 illustrates the flowchart of the proposed rendering-oriented point cloud compression framework. The framework is divided into two primary modules: the compression module and the rendering module. The compression module efficiently compresses the input point clouds to compact code streams and decompresses them to reconstructed point clouds. Following the convention of point cloud attribute compression methods [43, 21, 45, 22], we assume that the geometry has been losslessly transmitted. The rendering module renders the reconstructed point clouds to images from preset multiple perspectives through differentiable rendering techniques. Note that the rendering component of our framework is modular, allowing for the potential integration of future advancements in differentiable rendering techniques to replace the current rendering module used in this study.
@@ III-B Compression Module
where $\vec{f}_{i}^{\prime}$ is the output feature and $\delta_{\mathrm{rel}}\left(\vec{c}_{i}-\vec{c}_{j}\right)$ is the relative position encoding. Functions $\theta$, $\alpha$ and $\lambda$ are the multi-layer perceptron-based query, key and value projection layers, respectively. The local self-attention layer utilizes the cosine similarity to measure the relationship between the $\theta\left(\vec{f}_{i}\right)$ and $\alpha\left(\vec{f}_{j}\right)+\delta_{\mathrm{rel}}\left(\vec{c}_{i}-\vec{c}_{j}\right)$ vectors, rather than the commonly used $\operatorname{softmax}(\cdot)$ function. This choice stems from the adaptive nature of the fixed-size window. Firstly, by focusing on the direction, it inherently handles similarity and makes a fair comparison when the number of neighbors varies. In contrast, the similarity calculated from the $\operatorname{softmax}(\cdot)$ function highly depends on the number of neighbors. Secondly, it prevents the potential issue of the $\operatorname{softmax}(\cdot)$ function. A single neighbor (i.e., $|\mathcal{N}(i)|=1$) would make the attention mechanism degenerate to a linear layer by normalizing the weight to unity. Thus, using cosine similarity in the local self-attention layer of SP-Trans is a proper choice that aligns with the adaptive, fixed-size window approach to neighbor searching. It ensures equitable weighting of neighbors that reflects the local point distribution. Combined with a skip connection, the final SP-Trans is illustrated in Fig. 5.
@@ III-C Rendering Module
Once the points are in the camera view coordinate system, they need to be transformed into the normalized device coordinate (NDC) $X_{\text{ndc}}$. The specific type of camera (perspective or orthographic) will determine the projection matrix used for this transformation. The perspective camera, which more closely conforms to real scene shooting is chosen in our framework. The coordinates are first placed into a symmetric viewing frustum which represents the region camera shot, as shown in Fig. 7(a). Then the corresponding projective transformation [56] [^1] This projective transformation refers equation (41) in Section Perspective Projections of Chapter Three-Dimensional Viewing on Page 334 of reference [56]. Due to minor differences in program implementation (i.e., left handed coordinate system vs right handed coordinate system and the range of mapped z coordinates [0,1] vs [-1,1] for more accurate depth precision), there are also minor differences in the z coordinate of NDC coordinates. However, due to the screen coordinates calculated from x and y components of NDC coordinates and the invariance of z (i.e., depth)-orders of points, this does not influence the effect of rendering, as indicated by Eq. (9) and Eq. (10)., associated with a perspective division can be represented as
@@
$\tilde{X}_{\text{ndc}}$ $=\left[\begin{matrix}\frac{1}{\tan\left(\frac{fov}{2}\right)\times ap}&0&0&0\\ 0&\frac{1}{\tan\left(\frac{fov}{2}\right)}&0&0\\ 0&0&\frac{far}{far-near}&-\frac{far\times near}{far-near}\\ 0&0&1&0\end{matrix}\right]\tilde{X}_{\text{view}}$ (8)
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=\left[\begin{matrix}\frac{x}{\tan\left(\frac{fov}{2}\right)\times ap}\\ \frac{y}{\tan\left(\frac{fov}{2}\right)}\\ \frac{far\left(z-near\right)}{\left(far-near\right)}\\ z\end{matrix}\right]}\simeq\left[\begin{matrix}\frac{x}{z\times\tan\left(\frac{fov}{2}\right)\times ap}\\ \frac{y}{z\times\tan\left(\frac{fov}{2}\right)}\\ \frac{far\left(z-near\right)}{z\left(far-near\right)}\\ 1\end{matrix}\right],$
@@
where $fov$, $far$ and $near$ are the parameters of the viewing frustum. Specifically, $fov$ is the field of view angle, $near$ is the distance of the camera to the near clipping plane, $far$ is the distance to the far clipping plane and $ap$ is the aspect ratio of the image pixels. The transformation of points in camera view coordinate system to NDC coordinates need to be calculated in the perspective space, where the coordinates are represented as homogeneous coordinates. $\sim$ means the homogeneous coordinates of corresponding terms. $\left[x,y,z,1\right]^{\top}$ are the homogeneous coordinates of $\tilde{X}_{\text{view}}$. $\left[\frac{x}{\tan\left(\frac{fov}{2}\right)\times ap},\frac{y}{\tan\left(\frac{fov}{2}\right)},\frac{far\left(z-near\right)}{\left(far-near\right)},z\right]^{\top}$ are the homogeneous coordinates of $\tilde{X}_{\text{ndc}}$ after the projective transformation. Then perspective division actually transforms the viewing frustum in the camera system into the cubic space in the NDC system, as shown in Fig 7(b), which divides each component by the homogeneous $w$-component (i.e., $z$). Finally, we can get the 3D coordinates of $X_{ndc}$, i.e., $\small\left[\frac{x}{z\times\tan\left(\frac{fov}{2}\right)\times ap},\frac{y}{z\times\tan\left(\frac{fov}{2}\right)},\frac{far\left(z-near\right)}{z\times\left(far-near\right)}\right]^{\top}$. $\simeq$ denotes the equality under the homogeneous coordinates. In the subsequent and final viewport transformation, the $x$ and $y$ components will be transformed to the 2D screen coordinates by linear transformations. From the coordinates of ${X}_{\text{ndc}}$, we can see that the $x$ and $y$ components are scaled by the $z$-component (i.e., depth value). As a result, objects that are closer to the viewer appear larger than those that are farther away. This is how the projective transformation altered the geometry of the point cloud. As shown in Fig. 8, square grids of the chessboard will appear to converge and become trapezoidal in the distance. Without loss of generality, the farther an object is from the viewer, the more distorted its shape becomes.
Finally, the viewport transformation, determined by the parameters of the rendering window such as its resolution, maps the first two dimensions of the NDC coordinates, which range from $-1$ to $1$, to 2D screen coordinates. To provide a sense of depth and realism to the rendered 2D image, these sequential transformations present these points in 3D world with a new geometric relationship in 2D image.
@@
Subsequently, based on the altered geometry, the rasterization algorithm sorts the points in descending order of their depth values in NDC system to ensure that the closest points are processed first as shown in Fig. 9. This is achieved by calculating the depth value for each point and storing it in a z-buffer. Each point is splatted to a circular region in screen space, with opacity decreases away from the region’s center. The value of each pixel is computed by blending information from the K-nearest points in the z-buffer whose splatted regions overlap the pixel.
@@ IV-A1 Datasets
  • Training datasets. The model is trained on 548 point clouds sampled from 3D models in the real-world textured things (RWTT) dataset [65]. For each point cloud, $1\times 10^{6}$ points are randomly sampled from the surface and then voxelized to a $1024^{3}$ space to maintain consistency with the test datasets. Due to the GPU memory limitation, we partition these 3D point clouds into 3D patches. We show four training samples in Fig 10(a).
  • Testing datasets. Eight point clouds from two datasets used in MPEG are included in our testing samples. As shown in Fig. 10(b) and Fig. 10(c), four point clouds are defined in 8iVFB [32] and the other four are defined in Owlii [33]. ScanNet [66] is a large-scale indoor point cloud dataset containing 1603 scans at 2cm geometry precision. We used the officially divided training set of 1503 point clouds for training and the official test set of 100 point clouds for testing. Some samples in ScanNet are shown in Fig 10(d). The training and testing datasets are noticeably different, which demonstrates the generalization of our framework.
@@ IV-A2 Training and testing details
  • Training phase. The ground truth multiviews include 8 views with elevations of $0^{\circ}$ and azimuths from $0^{\circ}$ to $360^{\circ}$ with an interval of $60^{\circ}$ and 2 views with elevations of $90^{\circ}$ and $270^{\circ}$ and azimuth of $0^{\circ}$, as shown in Fig. 11. To maximize the coverage of the training patches in the rendered images to emulate the effect of using 1024x1024x1024 images for training, we correspondingly reduced the training image resolution empirically to 384x384. For the loss function, the value of $\lambda$ is set to 25,000, 4,000, 800, 250, and 85 for the training of the respective models. Each model undergoes training consisting of 140 epochs, facilitated by the PyTorch platform and the Adam optimizer. The parameters $\beta_{1}$ and $\beta_{2}$ of the Adam optimizer are configured at 0.9 and 0.999, respectively. The learning rate starts at 1e-4 and is progressively reduced to 1e-6, with a reduction to three-fourths of its previous value occurring every 15 epochs. The batch size is set to 8. The experimental setup includes a computer equipped with an Intel i9-14900k CPU operating at 3.60 GHz, 32 GB of main memory, and an NVIDIA GeForce RTX 4090 GPU.
  • Testing phase. To demonstrate the generalization of RO-PCAC to different views, the ground truth multiviews include 6 views with the elevation of $0^{\circ}$ and azimuths from $30^{\circ}$ to $360^{\circ}$ with an interval of $60^{\circ}$. The resolution of each view’s image is 1024$\times$1024, which matches the resolution of the test point cloud. The R-D performance includes bpp for rate, PSNR and MS-SSIM of rendered images for distortion. Furthermore, the BD-BR (Bjøntegaard Delta Rate) [67] is used to evaluate the overall compression efficiency. The R-D performance in both Y and YUV (6:1:1) spaces [68] is reported.